OpenStudy (anonymous):

Which of the following is the simplified form of fifth root of x times the fifth root of x times the fifth root of x times the fifth root of x? x to the 1 over fifth power x to the 4 over fifth power x to the four over twentieth power x

OpenStudy (anonymous):

im going with c

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

\[\sqrt[5]{x}\sqrt[5]{x}\sqrt[5]{x}\sqrt[5]{x}\] can be simplified further into x^4/5, so yes it's c.

OpenStudy (freckles):

\[x^\frac{1}{5} x^\frac{1}{5} x^\frac{1}{5} x^{\frac{1}{5}} =x^{\frac{1}{5}+\frac{1}{5}+\frac{1}{5}+\frac{1}{5}}\]